A balloon has a volume of 4.0 L when at sea level (1.0 atm) at room temperature 0f 28 . C What will be its volume when inflated with inflated with the same amount of gas at an elevation where the atmospheric pressure is 700 mm Hg at 28 . C


1
Expert's answer
2021-05-31T06:18:49-0400

The ideal-gas equation



"PV=nRT"


The amount of the gas is the same.

Then



"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}=\text{constant}"

The balloon has a volume of 4.0 L when at sea level (1.0 atm) at a room temperature of 28 °C.

"V_1=4.0\ L=4\times 10^{-3}\ m^3,"

"P_1=1.0\ atm=101325\ Pa,"

"T_1=28\degree C=(28+273)\ K=301\ K"


"1 mm\ Hg = 133.322\ Pa"

"P_2=700mm\ Hg=(700\cdot 133.322)\ Pa=93325.4\ Pa"

 "T_2=28\degree C=(28+273)\ K=301\ K=T_1"          




"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"

Solve for "V_2"



"V_2=\dfrac{P_1T_1}{P_2T_2}V_1"

Substitute

    


"V_2=\dfrac{101325\ Pa(301\ K)}{93325.4\ Pa(301\ K)}\cdot4\times 10^{-3}\ m^3""\approx 4.343\times 10^{-3}\ m^3\approx4.343\ L"

The volume of gas will be "V_2=4.343\ L."